# Enhancing Score-Based Sampling Methods with Ensembles

### v1

@@ 2.4.2 Antithetic sampling
$\hat{p}_{t}(\boldsymbol{x})$ $=\frac{\sum_{i=1}^{N_{\rm{ens}}}\kappa_{t}(\boldsymbol{x}|\boldsymbol{x}^{\prime}_{i})\frac{p_{0}(\boldsymbol{x}^{\prime}_{i})}{p_{\rm{is}}(\boldsymbol{x}^{\prime}_{i})}}{2N_{\rm{ens}}}$ (9)
$+\frac{\sum_{i=1}^{N_{\rm{ens}}}\kappa_{t}(\boldsymbol{x}|-\boldsymbol{x}^{\prime}_{i})\frac{p_{0}(-\boldsymbol{x}^{\prime}_{i})}{p_{\rm{is}}(-\boldsymbol{x}^{\prime}_{i})}}{2N_{\rm{ens}}}.$
This strategy is a common variance reduction strategy in Bayesian contexts (e.g., Geweke, 1988). The impact on the quality of the score estimator is shown in Figure 2.
@@ 2.4.3 Periodic Resampling Strategy
Since the variance of the estimator $\hat{p}_{t}$ depends on $p_{\rm{is}}$, which itself depends on the statistics of the ensemble $\{\boldsymbol{x}_{i}\}$, we can refine $\hat{p}_{t}$ during the reverse diffusion by updating $p_{\rm{is}}$ using the statistics of $\{\boldsymbol{x}_{i}\}$ at discrete resampling times, $t_{r}$, where $10<t_{\rm{r}}<$ 1. After each resampling time, $\hat{p}_{t}$ is held fixed until the next resampling time. By increasing the number of resampling times, $N_{\rm{r}}$, during reverse diffusion, $p_{\rm{is}}$ is adapted to remain in the proximity of $p_{0}$ during the diffusion process, which prevents the likelihood ratio $p_{0}(\boldsymbol{x}^{\prime})/p_{\rm{is}}(\boldsymbol{x}^{\prime})$ from getting too small to effectively perform importance sampling (Agapiou et al., 2017). Importantly, $p_{0}$ is only evaluated when updating $\hat{p}_{t}$, so there is a trade-off between increasing $N_{\rm{r}}$ to improve the score estimate while decreasing $N_{r}$ to limit the total number of evaluations of $p_{0}$. For the experiments shown here, $N_{r}$ is generally set to between 10 and 30. Overall, this strategy is analogous in spirit to MCMC methods that utilize proposal distributions that adapt to the statistics of the Markov chains during sampling (e.g., Haario et al., 2001; Minson et al., 2013). The outline of this process is shown in Algorithm 1.
@@
By recomputing $p_{\rm{is}}$ only periodically, we reduce the number of $p_{0}$ evaluations, which is beneficial when $p_{0}$ is expensive to evaluate. This situation is common in complex black box simulations settings where not only gradients are unavailable, but evaluations of the black box are also computationally expensive. As a result, we limit the number of forward evaluations in Algorithm 1 to a total of $N_{\rm{ens}}N_{\rm{r}}$. For example, $N_{\rm{r}}=10$ resampling steps and an ensemble size $N_{\rm{ens}}=100$ would result in $1000$ evaluations of $p_{0}$.
@@ 2.5 Comparison with Ensemble Kalman Methods
Sampling and optimization methods based on Ensemble Kalman methods use an ensemble covariance matrix to estimate the gradient of a log probability density function that serves as the drift within a Langevin-type stochastic differential equation, see for example Garbuno-Inigo et al. (2020) for an easy-to-digest exposition. Usually, one can show that these methods become exact samplers in a Gaussian setting and convergence results can be obtained, for example, in convex and mean field settings (Schillings & Stuart, 2017; Garbuno-Inigo et al., 2020; Calvello et al., 2022). Score-based reverse diffusion samplers, as presented here, have the added advantage that they become exact samplers for nearly any target distribution as the ensemble size and resampling times tend to infinity. However, score-based methods, as presented here, have the added disadvantage that they fight an uphill battle against the curse of dimensionality, discussed further in Sections 3.2 and 3.3. This situation is not surprising as estimating the score function in high dimensions using Monte Carlo sampling alone is a losing proposition because controlling the variance of the estimator becomes increasingly difficult.
@@ 3 Experiments
3.1 Low-dimensional Problems
We first apply our ensemble strategy to a set of two-dimensional Bayesian toy problems that exhibit posterior distributions with varying degrees of multimodality and non-ellipticity, which generally provide challenges to many MCMC techniques for sampling probability distributions. As shown in Figure 1, these problems include the well-known “banana”-shaped distribution, a distribution with neighboring high-probability ridges (adapted from Morzfeld et al. (2018)), and a mixture of three Gaussian distributions with well-separated means. The parameters used for each experiment are specified in Table 1.
Figure 1: Comparison of different methods for drawing random samples from three different two-dimensional probability distributions. For each distribution, the joint plot shows contours of log probability along with the random samples while the marginal plots show kernel density estimates of the marginal distributions. The leftmost column shows direct samples from the distributions, while the rest of the columns represent samples from a variational inference approximation (VI), the No U-Turn Sampler (NUTS), a Metropolis-adjusted Langevin Algorithm (MALA), and the our approach (ENS).
Table 1: Parameters for the examples shown in Figure 1. For these experiments, the forward process has zero drift (i.e., $\boldsymbol{b}_{t}=\boldsymbol{\mu}=0$), and the diffusion $\boldsymbol{g}_{t}=(\sigma_{\rm{min}}^{1/p}+t(\sigma_{\rm{max}}^{1/p}-\sigma_{\rm{min}}^{1/p}))^{p}$, where $p=5$ (Karras et al., 2022). MIS here stands for the multiple importance sampling approach from Appendix E. Additionally, $N_{\rm{ens}}=1000$, $N_{r}=10$, and the initial time step size is specified by $\Delta t_{\rm{init}}$.
@@ 3.1 Low-dimensional Problems
For comparison, we generate samples from these target distributions using three different techniques: 1) approximation of the distribution with a 2D multivariate Gaussian with parameters estimated via minimization of the reverse KL-divergence (i.e., variational inference (VI)); 2) NUTS; and 3) MALA.
As expected, the non-ellipticity of all three distributions present difficulties for methods that rely on or are related to Gaussian approximations. The reverse KL-divergence cost function used for VI encourages mode-seeking approximations, which can miss substantial areas of high probability for skewed or multi-modal distributions (Blei et al., 2017). Alternatively, both NUTS and MALA perform well for the “banana” distribution, but NUTS has a tendency to draw samples from a single mode or ridge of high probability, even when the ridges are not well-separated. MALA is most similar to our method in that a diffusion process controls the random walk of samples that are initially randomly distributed, which permits samples to reach multiple regions of high probability. However, Langevin dynamics can suffer from slow mixing when modes are separated by low density regions, causing MALA to generate samples with incorrect relative densities for the mixture of Gaussians distribution (Song & Ermon, 2019). Our method is able to correctly sample all three distributions, demonstrating its ability to handle both non-ellipticity and multi-modality. We note that while a simple Metropolis-Hastings sampler could also accurately sample from these 2D distributions (with a large enough variance for the proposal distribution), it would quickly lose efficiency in higher dimensions. Our results broadly confirm the results obtained in Huang et al. (2021) where a Föllmer sampler was used to generated results from two-dimensional distributions, albeit their approach differs from ours in several ways (e.g., different Monte Carlo estimators, number of Monte Carlo samples, etc.)
@@ 3.1.1 Influence of ensemble size
As shown in Equations (6) and (2.4.1), the ensemble size affects the accuracy of the Monte Carlo estimate for the score function. Incorrect drift estimates will cause the sample trajectories to move toward incorrect locations in parameter space. In practice, the minimum ensemble size for sufficiently accurate drift estimates will depend on the dimensionality and “complexity” of the probability distribution. To explore this effect, we use our method to sample from a five-dimensional multivariate Gaussian distribution for different ensemble sizes. We quantify sample accuracy on the final ensembles generated by our method with an energy-based distance measurement that estimates the statistical distance between two sets of random vectors (Appendix B). We find that the energy distance of our method (measured between generated and reference samples from the 5-$d$ Gaussian) is reduced by a factor of three when increasing $N_{\rm{ens}}$ from 16 to 64, with a further factor of two reduction with $N_{\rm{ens}}$ from 64 to 256 (Figure 2). For smaller ensemble sizes, the antithetic sampling strategy can significantly decrease sampling errors, while for larger ensembles, antithetic sampling can minimize the occurrence of high energy distances associated with inaccurate diffusion trajectories.
@@
Figure 2: Violin plots showing energy distance between samples generated by our methods and ground truth samples from a 5-$d$ Gaussian as a function of ensemble size and dependent on whether an antithetic sampling strategy is used. At each ensemble size, energy distance was computed for $200$ random permutations of the samples. The violin plots are bounded by the 10th and 90th percentiles of the resulting energy statistics.

### v2

@@ 3.2.2 Antithetic sampling
$\hat{p}_{t}(\boldsymbol{x})$ $=\frac{\sum_{i=1}^{N_{\rm{ens}}}\kappa_{t}(\boldsymbol{x}|\boldsymbol{x}^{\prime}_{i})\frac{p_{0}(\boldsymbol{x}^{\prime}_{i})}{p_{\rm{is}}(\boldsymbol{x}^{\prime}_{i})}}{2N_{\rm{ens}}}+\frac{\sum_{i=1}^{N_{\rm{ens}}}\kappa_{t}(\boldsymbol{x}|-\boldsymbol{x}^{\prime}_{i})\frac{p_{0}(-\boldsymbol{x}^{\prime}_{i})}{p_{\rm{is}}(-\boldsymbol{x}^{\prime}_{i})}}{2N_{\rm{ens}}}.$ (9)
This variance reduction strategy is common in Bayesian contexts (e.g., [25]). The impact on the quality of the score estimator is shown in Figure 4.
@@ 3.2.3 Periodic Resampling Strategy
Since the variance of the estimator $\hat{p}_{t}$ depends on $p_{\rm{is}}$, which itself depends on the statistics of the ensemble $\{\boldsymbol{x}_{i}\}$, we can refine $\hat{p}_{t}$ during the reverse diffusion by updating $p_{\rm{is}}$ using the statistics of $\{\boldsymbol{x}_{i}\}$ at discrete resampling times, $t_{r}$, where $10<t_{\rm{r}}<$ 1. After each resampling time, $\hat{p}_{t}$ is held fixed until the next resampling time. By increasing the number of resampling times, $N_{\rm{r}}$, during reverse diffusion, $p_{\rm{is}}$ is adapted to remain in the proximity of $p_{0}$ during the diffusion process, which prevents the likelihood ratio $p_{0}(\boldsymbol{x}^{\prime})/p_{\rm{is}}(\boldsymbol{x}^{\prime})$ from getting too small to effectively perform importance sampling [26]. Importantly, $p_{0}$ is only evaluated when updating $\hat{p}_{t}$, so there is a trade-off between increasing $N_{\rm{r}}$ to improve the score estimate while decreasing $N_{r}$ to limit the total number of evaluations of $p_{0}$. For the experiments shown here, $N_{r}$ is generally set to between 10 and 30. Overall, this strategy is analogous in spirit to MCMC methods that utilize proposal distributions that adapt to the statistics of the Markov chains during sampling (e.g., [27, 28]). The outline of this process is shown in Algorithm 1.
@@
By recomputing $p_{\rm{is}}$ only periodically, we reduce the number of $p_{0}$ evaluations, which is beneficial when $p_{0}$ is expensive to evaluate. This situation is common in complex black box simulations settings where not only gradients are unavailable, but evaluations of the black box are also computationally expensive. With this approach, the total number of forward evaluations in Algorithm 1 is limited to $N_{\rm{ens}}N_{\rm{r}}$. For example, $N_{\rm{r}}=10$ resampling steps and an ensemble size $N_{\rm{ens}}=100$ would result in $1000$ evaluations of $p_{0}$.
3.2.4 Anisotropic Forward Diffusion
For certain classes of problems where some information about the covariance structure of the parameters is available a priori, we can employ a localization-type strategy. Specifically, the flexible form of Equation (1) allows us to choose the scaling matrix of the forward diffusion to be a localization matrix defined as the Cholesky-decomposed prior covariance. In this way, we can “push” the diffusion process to stay “close” to the dominant subspace defined by the prior covariance. The implicit assumption is that the posterior will follow a similar covariance structure as defined by the prior. In essence, this localization strategy is a kind of linear dimensionality reduction strategy where we assume that we know the relevant subspace a priori (e.g., [29]). As a result, this strategy becomes more useful as the dimensionality of the problem increases. In the examples presented here, we utilize an Ornstein-Uhlenbeck forward process for the forward diffusion process with drift and diffusion operators defined as
$\boldsymbol{b}_{t}$ $=\theta(\boldsymbol{\mu}-\boldsymbol{x}_{t}),$ (10a)
which evolves samples from $p_{0}$ to a multivariate Gaussian distribution with mean $\boldsymbol{\mu}$ and covariance matrix $\alpha\boldsymbol{\Sigma_{\text{prior}}}$ over a time scale $\theta>0$. Here, $\alpha$ is a scalar inflation factor for scaling the prior covariance in order to optionally inflate the initial ensemble. We choose the magnitude of $\theta$ such that as $t\rightarrow 1$ the process has converged to $\mathcal{N}(\boldsymbol{\mu},\alpha\boldsymbol{\Sigma}_{\text{prior}})$ in a distributional sense. The reverse diffusion process can then be determined by simply plugging $\boldsymbol{b}_{t}$ and $\boldsymbol{g}_{t}$ into Equation (2).
The utility of the Ornstein-Uhlenbeck forward process can be illustrated with a simple 2D example of forward diffusion from a multivariate normal distribution. For scalar Brownian motion, the samples will take random walks in each dimension with equal variance. For higher-dimensional problems, this behavior can cause many samples to move towards low probability regions of the target distribution. With the Ornstein-Uhlenbeck process, the random walks are constrained to the covariance structure imposed by $\boldsymbol{g}_{t}$, which for certain problems can keep samples in the relevant regions of the probability space to improve the estimate of the Föllmer drift (Figure 2). In practice, we also tune $\theta$ to control the relative importance of $\mu$ during the reverse diffusion. For $\theta=1$, the samples will stay close to $\mu$, which can be un-desirable when the target mean is expected to differ non-negligibly from $\mu$. For the higher-dimensional experiments in this work, we find $\theta=0.1-0.2$ to provide the best accuracy.
@@ 3.3 Comparison with Ensemble Kalman Methods
Sampling and optimization methods based on Ensemble Kalman methods use an ensemble covariance matrix to estimate the gradient of a log probability density function that serves as the drift within a Langevin-type SDE, see for example [17] for an easy-to-digest exposition. Usually, one can show that these methods become exact samplers in a Gaussian setting and convergence results can be obtained, for example, in convex and mean field settings [9, 16, 17]. Score-based reverse diffusion samplers, as presented here, have the added advantage that they become exact samplers for nearly any target distribution as the ensemble size and resampling times tend to infinity (see Section 4.2 for a numerical comparison of the two methods for the Lorenz63 system). However, score-based methods, as presented here, have the added disadvantage that they fight an uphill battle against the curse of dimensionality, discussed further in Sections 4.3 and 4.4. This situation is not surprising as estimating the score function in high dimensions using Monte Carlo sampling alone is a losing proposition because controlling the variance of the estimator becomes increasingly difficult. In short, we expect similar sampling accuracy between our method and Ensemble Kalman methods for near-Gaussian probability distributions and improved accuracy for our method for non-Gaussian distributions.
@@ 4 Experiments
In this section, we demonstrate the effectiveness of our proposed ensemble strategies for score-based sampling by conducting a series of comprehensive tests across a variety of sampling problems, ranging from low- to high-dimensional scenarios. We provide a comparison against performant gradient-based MCMC techniques and demonstrate improvements in sampling accuracy without the need for gradient information.
4.1 Two-dimensional Problems
We first apply our ensemble strategy to a set of two-dimensional Bayesian toy problems that exhibit posterior distributions with varying degrees of multimodality and non-ellipticity, which generally provide challenges to many MCMC techniques for sampling probability distributions. As shown in Figure 3, these problems include the well-known “banana”-shaped distribution, a distribution with neighboring high-probability ridges (adapted from [30]), and a mixture of three Gaussian distributions with well-separated means. The parameters used for each experiment are specified in Table 1.
Figure 3: Comparison of different methods for drawing random samples from three different two-dimensional probability distributions. For each distribution, the joint plot shows contours of log probability along with the random samples while the marginal plots show kernel density estimates of the marginal distributions. The leftmost column shows direct samples from the distributions, while the rest of the columns represent samples from a variational inference approximation (VI), the No U-Turn Sampler (NUTS), a Metropolis-adjusted Langevin Algorithm (MALA), and our approach (ENS).
Table 1: Parameters for the examples shown in Figure 3. For these experiments, the forward process has zero drift (i.e., $\boldsymbol{b}_{t}=\boldsymbol{\mu}=0$), and the diffusion $\boldsymbol{g}_{t}=(\sigma_{\rm{min}}^{1/p}+t(\sigma_{\rm{max}}^{1/p}-\sigma_{\rm{min}}^{1/p}))^{p}$, where $p=5$ [31]. MIS here stands for the multiple importance sampling approach from Appendix B. Additionally, $N_{\rm{ens}}=1000$, $N_{r}=10$, and the initial time step size is specified by $\Delta t_{\rm{init}}$.
@@ 4.1 Two-dimensional Problems
For comparison, we generate samples from these target distributions using three different techniques: 1) approximation of the distribution with a 2D multivariate Gaussian with parameters estimated via minimization of the reverse KL-divergence (i.e., variational inference (VI)); 2) NUTS; and 3) MALA. Note that for NUTS, a single chain is used for sampling, whereas for MALA, multiple chains are initialized with random states and only the final states are used as samples of the target distribution. In D, we also show the performance of two established gradient-free sampling algorithms: i) Population Monte Carlo (PMC), and ii) Ensemble Slice Sampling (ESS).
Method Banana Ridged Mixture
VI 39.280 $\pm$ 9.100 6.698 $\pm$ 2.033 14.822 $\pm$ 5.038
NUTS 3.981 $\pm$ 3.360 20.435 $\pm$ 3.032 16.620 $\pm$ 3.682
MALA 7.143 $\pm$ 4.894 0.891 $\pm$ 0.672 6.768 $\pm$ 3.397
ESS 4.436 $\pm$ 3.127 0.592 $\pm$ 0.437 16.623 $\pm$ 3.667
PMC 2.969 $\pm$ 2.521 0.504 $\pm$ 0.444 9.352 $\pm$ 3.794
ENS 3.102 $\pm$ 2.616 0.531 $\pm$ 0.389 1.945 $\pm$ 1.835
Table 2: Comparison of energy distance (mean and standard deviation) for the two-dimensional experiments using different sampling methods.
As expected, the non-ellipticity of all three distributions present difficulties for methods that rely on or are related to Gaussian approximations (which includes the aforementioned Ensemble Kalman methods). Moreover, the reverse KL-divergence cost function used for VI encourages mode-seeking approximations, which can miss substantial areas of high probability for skewed or multi-modal distributions [32]. Alternatively, both NUTS and MALA perform well for the “banana” distribution, but for the second target distribution, NUTS with a single chain has a tendency to draw samples from a single mode or ridge of high probability, even when the ridges are not well-separated. Multiple runs of NUTS with different random initializations would likely improve sampler exploration in this case. MALA is most similar to our method in that a diffusion process controls the random walk of samples that are initially randomly distributed, which permits samples to reach multiple regions of high probability. However, Langevin dynamics can still suffer from slow mixing when modes are separated by low density regions (as is the case for the third distribution), causing MALA to generate samples with incorrect relative densities for the mixture of Gaussians distribution [22].
To quantify sampling accuracy, an energy-based distance measurement (A) that estimates the statistical distance between two sets of random vectors is computed for the samples relative to samples from the true distribution (Table 2). Our method is able to correctly sample all three distributions, demonstrating its ability to handle both non-ellipticity and multi-modality. These results broadly confirm the results obtained in [3] where a Föllmer sampler was used to generated results from two-dimensional distributions, albeit their approach differs from ours in several ways (e.g., different Monte Carlo estimators, number of Monte Carlo samples, etc.)
@@ 4.1.1 Influence of Ensemble Size
As shown in Equations (6) and (7), the ensemble size affects the accuracy of the Monte Carlo estimate for the score function. Incorrect drift estimates will cause the sample trajectories to move toward incorrect locations in parameter space. In practice, the minimum ensemble size for sufficiently accurate drift estimates will depend on a multitude of factors, including the dimensionality of the probability distribution, the choice of importance sampling distribution, and the design of the forward process (C). We follow the practice of Ensemble Kalman methods by starting with an ensemble size that is 10 times the dimensionality of the probability distribution [e.g., 16, 17]. We then compare the final posterior samples with those obtained from an ensemble size that is 1.5–2 times larger. If the samples exhibit substantially different marginal distributions, then we increase the ensemble size further.
@@
To explore this effect, we use our method to sample from a five-dimensional multivariate Gaussian distribution for different ensemble sizes and measure the energy distance relative to the reference 5-$d$ Gaussian. We find that the energy distance of our method is reduced by a factor of three when increasing $N_{\rm{ens}}$ from 16 to 64, with a further factor of two reductions with $N_{\rm{ens}}$ from 64 to 256 (Figure 4). These reductions scale roughly with $\sqrt{N_{\rm{ens}}}$, consistent with the convergence analysis in C. For smaller ensemble sizes, the antithetic sampling strategy can significantly decrease sampling errors, while for larger ensembles, antithetic sampling can minimize the occurrence of high energy distances associated with inaccurate diffusion trajectories. These trends are broadly consistent across different target probability distributions, but the exact ensemble size at which accuracy will degrade will require experimentation on a case-by-case basis. More robust and automated selection of ensemble sizes remains an active area of research.
@@
Figure 4: Violin plots showing energy distance between samples generated by our methods and ground truth samples from a 5-$d$ Gaussian as a function of ensemble size, $N_{\rm{ens}}$, and dependent on whether an antithetic sampling strategy is used. At each ensemble size, energy distance was computed for $200$ random permutations of the samples. The violin plots are bounded by the 10th and 90th percentiles of the resulting energy statistics. Black and red dashed lines are the theoretical $1/\sqrt{N_{\rm{ens}}}$ and $1/N_{\rm{ens}}$ curves, respectively.
4.1.2 Influence of Ensemble Initialization
Similar to the ensemble size, the seeding of the initial ensemble members will affect the accuracy of the estimated score function. Generally, the Monte Carlo estimate requires that a few ensemble members have sufficiently high initial likelihood values in order to estimate an accurate score using Equation (6). We can see this effect in Figure 5 where we sample from the 2D banana distribution with three different initial ensembles.
Figure 5: Reverse diffusion trajectories for different initial ensembles for the banana probability distribution. The diffusion times span from $t=1$ (left) to $t=0$ (right), and the different rows correspond to different initial ensembles with varying offsets relative to the true distribution. Black dashed elliptical contours correspond to the importance sampling distribution, $p_{\text{is}}$, at the given diffusion times. In these examples, a Gaussian $p_{\text{is}}$ is used. Overall, ensembles with larger offsets will lead to undersampling of the distribution.
If only a few ensemble members have locations that lie close to the target probability distribution, the rest of the ensemble will evolve towards those locations, leading to undersampling of the distribution. However, even when the initial ensemble lies entirely outside of the distribution (e.g., the last row in Figure 5), the variance of the importance sampling distribution can still introduce a sufficient dynamic range in likelihood values to allow the ensemble to evolve in the correct direction. Therefore, an iterative approach may be applied where the reverse diffusion ensemble is re-initialized with a bias determined from the mean of the ensembles from previous runs. In this way, the initial ensemble has greater overlap with the target distribution.
4.2 Three-dimensional Lorenz63 Problem
The Lorenz63 system is a well-known simplified model for atmospheric convection consisting of the following three ordinary differential equations:
$\frac{dy_{1}}{dt}$ $=\sigma(y_{2}-y_{1}),$ (11)
$\frac{dy_{2}}{dt}$ $=y_{1}(r-y_{3})-y_{2},$ (12)
$\frac{dy_{3}}{dt}$ $=y_{1}y_{2}-\beta y_{3},$ (13)
where $\sigma$, $r$, and $\beta$ are scalar system parameters that are normally assumed to be positive. The trajectories of $y_{1}$, $y_{2}$, and $y_{3}$ can be computed with a standard ODE solver with initial conditions sampled from a unit normal distribution. For certain combinations of $\sigma$, $r$, and $\beta$, the trajectories will exhibit chaotic behavior that is highly sensitive to the initial conditions. This behavior prevents the reliable computation of gradients of the system outputs with respect to system parameters, which ultimately prevents the use of gradient-based MCMC methods to estimate the parameter values from observations. Here, we use the values $\sigma=10$, $r=24.5$, and $\beta=8/3$, which places the system near the transition between stable fixed points and chaotic behavior.
[…]
4.2.1 Sampling the Posterior Distribution
The Bayesian inference problem formulates the un-normalized posterior distribution for $\mathbf{x}$ as:
where the first distribution on the right-hand side is the likelihood distribution and the second distribution is the prior distribution. Here, we use a multivariate normal distribution for the likelihood with the mean vector and covariance matrix as described previously. For the prior, we use an independent normal distribution, $p(\boldsymbol{x})=\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\sigma^{2}I})$ with $\boldsymbol{\mu}=[12,25,8/3]$ and $\boldsymbol{\sigma}=[2,1,1]$. In practical applications, since $\mathbf{x}$ is strictly positive, we should sample for $\log(\boldsymbol{x})$, which would lead to a near-Gaussian posterior $p(\log(\boldsymbol{x})|\boldsymbol{d})$. Since one of our goals for this experiment is to highlight the difference in sampling behavior between our proposed method and Ensemble Kalman Sampling (EKS) [17], we opt to sample for $\mathbf{x}$, which will lead to non-Gaussian behavior for $p(\boldsymbol{x}|\boldsymbol{d})$. Furthermore, we multiply the covariance matrix by a factor of four when parameterizing the likelihood distribution in order to enhance the effect of the non-linear forward model on $p(\boldsymbol{x}|\boldsymbol{d})$.
For sampling $p(\boldsymbol{x}|\boldsymbol{d})$ with the reverse diffusion process, we prescribe a zero-drift forward diffusion process with $\boldsymbol{g}_{t}$ as in Table 1, $\sigma_{\text{min}}=0.05$, $\sigma_{\text{max}}=1$, $\Delta t_{\text{init}}=0.0025$, an MIS $p_{\text{is}}$, 20 resampling steps, and an ensemble size of $N_{\text{ens}}=1000$. For EKS, in order to prescribe the same number of forward model evaluations as our method, we use an ensemble size of 1000 for 20 iterations. Additionally, we sample $p(\boldsymbol{x}|\boldsymbol{d})$ with Population Monte Carlo (PMC), which is well suited for low-dimensional, non-Gaussian distributions.
The posterior samples derived from all three methods successfully encompass the true values (Figure 6). For the parameters $r$ and $\beta$, the marginal posterior is close to Gaussian, so the samples derived from all methods are nearly identical. However, the marginal posterior for $\sigma$ has a longer, low-probability tail for higher values. Since EKS utilizes Gaussian approximations for the posterior, the associated marginal samples for $\sigma$ have an inflated variance in order to span the tail, leading to oversampling of that region. In contrast, our score-based sampling method and PMC generate samples closer to the true posterior distribution. These results indicate that our proposed method is effective for Bayesian inference problems for highly non-linear systems that are infeasible to solve with gradient-based techniques.
Figure 6: Marginal posterior kernel density estimates for the three parameters ($\sigma$, $r$, $\beta$) of the Lorenz63 system. 1D marginals are computed from samples obtained with the Ensemble Kalman sampler (EKS; blue), Population Monte Carlo (PMC; orange), and the ensemble methods presented here (ENS; green). The true posterior probabilities are shown in black. Energy distances relative to the true posterior are shown in the bottom right plot. Overall, all methods perform well for sampling parameters that exhibit near-Gaussian marginal posterior distributions, while non-Gaussian behavior is more accurately sampled by the more flexible PMC and ensemble methods presented here.